Stan and Hilda can mow the lawn in 40 minutes if they work together. If Hilda works twice as fast as Stan, how long would it take Stan to mow the lawn alone?
Answer by stanbon(75887)

Stan and Hilda can mow the lawn in 40 minutes if they work together. If Hilda works twice as fast as Stan, how long would it take Stan to mow the lawn alone?
------
Stan time = 2x hr/job ; rate = 1/(2x) job /hr
---
Hilda time = x hr/job ; rate = 1/x job/hr
---
Together time = (2/3) hr/job ; rate = (3/2)job/hr
===
Equation:
rate + rate = together rate
1/x + 1/(2x) = 3/2
---
Multiply thru by 2x:
2 + 1 = 3x
---
x = 1 hr
Stan's time = 2 hr to do the job alone
=====
Cheers,
Stan H.